# strong corks derived from the Akbulut cork

### v1

@@ Abstract.
Using instanton theory, we prove that the boundaries of the corks introduced by Auckly, Kim, Melvin, and Ruberman in [AKMR14] are strong corks, and that any nontrivial linear combination of them also yields a strong cork. Furthermore, via Heegaard Floer theory, we show that they are strong corks with respect to a distinct involution, and we construct a larger family of strong corks that generalizes them. We also prove that the boundaries of the corks introduced by Tange [Tan16] are strong corks and that any nontrivial linear combination of them yields a strong cork.
@@ 1. Introduction
In [Akb91], Akbulut provided the first example of a corkcork, which is defined as a compact, contractible 4-manifold $W$ equipped with an orientation-preserving diffeomorphism $\tau$ on its boundary that does not extend to a diffeomorphism of WW. [^1] Some authors require $W$ to be Stein, such as in [AY08]. By Freedman’s theorem [Fre82], $\tau$ extends over $W$ as a homeomorphism. Akbulut proved that the smooth structure of $K3\#\overline{\mathbb{C}P^{2}}$ can be changed by cutting out an embedded copy of $W$ and regluing it via $\tau$. Such an operation is called a corktwistcork\ twist. More generally, Matveyev [Mat95] and Curtis-Freedman-Hsiang-Stong [CFcHS96] showed that any two exotic smooth structures on a closed, simply-connected 4-manifold are related by a cork twist. Numerous examples of corks have been constructed; see, for example, [AY08] and [Gom17]. Corks are typically detected by embedding them into a closed 4-manifold with a non-vanishing smooth 4-manifold invariant, or by utilizing Stein structures and the adjunction inequality [AM97].
@@
    The corks $(C_{n},\tau)$ and $(\overline{C}_{n},\tau)$, displayed in Figure 1, for even integers $n\geq 2$. These families were introduced by Auckly-Kim-Melvin-Ruberman in [AKMR14, Figure 8]. Note that $(C_{1},\tau)$ corresponds to the Akbulut cork $(W_{1},\tau)$, and $(\overline{C}_{1},\tau)$ corresponds to the first positron $(\overline{W}_{1},\tau)$. For odd $n$, it is shown in [DHM20] that $(\partial C_{n},\tau)$ and $(\partial\overline{C}_{n},\tau)$ are strong corks. The family of corks $(C_{n},\tau)$ was used to produce infinitely many exotic pairs of 4-manifolds in [Yas15].
@@ 1.1. Main results
We prove Theorems 1.2, 1.3, 1.4, and 1.5 in Section 3, and we prove Proposition 1.6 in Section 4 (appendix).
@@ 3.1. Heegaard-Floer-theoretic methods
$h_{\iota\circ\tau}(Y)=[(CF^{-}(Y)[-2],\iota\circ\tau)]\in\mathfrak{I}.$
The local equivalence group $\mathfrak{I}$ admits a partial order; see [DHM20, Definition 3.6]. The invariants $h_{\tau}(Y)$ and $h_{\iota\circ\tau}(Y)$ are monotonic under an appropriate equivariant negative-definite cobordism, as in the following theorem:
@@
We construct an interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)=(\Sigma(2,2m+1,4m+3),\sigma)$ as shown in Figure 13.222It can be verified that $Y_{m,1}$ is diffeomorphic to the boundary $\partial W_{m}$, where $(W_{m},\tau)$ is the Akbulut-Yasui cork in [AY08]. In [DHM20, Theorem 1.13], it is shown that $h_{\tau}(\partial W_{m})\leq h_{\tau}(B_{m})<0$, which implies that $(\partial W_{m},\tau)$ is a strong cork. This is formed by attaching (1)(-1)-framed 2-handles to the green unknots as in the top left of Figure 13. We see that the condition of Theorem 3.2 (2) are satisfied by simultaneously sliding the left green unknot over the right component of $Y_{m,1}$ and the right green unknot over the left component of $Y_{m,1}$. By Theorem 3.2 (2) and Lemma 3.3,
@@
Figure 13. The interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)$ used in the proof of Theorem 1.3
@@
Figure 14. The spinc-conjugating (1)(-1)-cobordism from $(Y_{m,n+1},\sigma)$ to $(Y_{m,n},\sigma)$ used in the proof of Theorem 1.3
@@ 3.3. Instanton-theoretic methods
[ADMT23, Theorem 7.7] Let $\{(Y_{i},\tau_{i})\}^{\infty}_{i=1}$ be a sequence of oriented integer homology 3-spheres equipped with orientation-preserving involutions $\tau_{i}$. Assume the fixed-point set of each $\tau_{i}$ is a copy of $S^{1}$, so that the equivariant connected sum operation is well-defined. Suppose that:
@@
Figure 15. The equivariant negative-definite cobordism $W_{m,n}$ from $(Z_{m,n+1},\tau)$ to $(Z_{m,n},\tau)$ used in the proof of Theorem 1.2.
@@
Figure 19. The simply-connected, equivariant negative-definite cobordism from $(S^{3}_{1/{m+1}}(K_{n}),\tau)$ to $(S^{3}_{1/m}(K_{n}),\tau)$ used in the proof of Theorem 1.5.

### v2

@@ Abstract.
We prove that the boundaries of the corks introduced by Auckly, Kim, Melvin, and Ruberman in [AKMR14] and by Tange in [Tan16] are strong corks. Furthermore, we prove that any nontrivial linear combination of them yields a strong cork, and we construct a larger family of strong corks that generalizes them. These results rely on the instanton-theoretic invariant introduced by Alfieri, Dai, Mallick, and Taniguchi in [ADMT23].
@@ 1. Introduction
In [Akb91], Akbulut provided the first example of a corkcork, which is defined as a compact, contractible 4-manifold $W$ equipped with an orientation-preserving diffeomorphism $\tau$ on its boundary that does not extend to a diffeomorphism of $W$. (Some authors require $W$ to be Stein, such as in [AY08].) By Freedman’s theorem [Fre82], $\tau$ extends over $W$ as a homeomorphism. Akbulut proved that the smooth structure of $K3\#\overline{\mathbb{C}P^{2}}$ can be changed by cutting out an embedded copy of $W$ and regluing it via $\tau$. Such an operation is called a corktwistcork\ twist. More generally, Matveyev [Mat95] and Curtis-Freedman-Hsiang-Stong [CFcHS96] showed that any two exotic smooth structures on a closed, simply-connected 4-manifold are related by a cork twist. Numerous examples of corks have been constructed; see, for example, [AY08] and [Gom17]. Corks are typically detected by embedding them into a closed 4-manifold with a non-vanishing smooth 4-manifold invariant, or by utilizing Stein structures and the adjunction inequality [AM97].
@@
    The corks $(C_{n},\tau)$ and $(\overline{C}_{n},\tau)$, displayed in Figure 1, for even integers $n\geq 2$. These families were introduced by Auckly-Kim-Melvin-Ruberman in [AKMR14, Figure 8]. (For [AKMR14], see also [Akb15].) Note that $(C_{1},\tau)$ corresponds to the Akbulut cork $(W_{1},\tau)$, and $(\overline{C}_{1},\tau)$ corresponds to the first positron $(\overline{W}_{1},\tau)$. For odd $n$, it is shown in [DHM20] that $(\partial C_{n},\tau)$ and $(\partial\overline{C}_{n},\tau)$ are strong corks. The family of corks $(C_{n},\tau)$ was used to produce infinitely many exotic pairs of 4-manifolds in [Yas15].
@@ 1.1. Main results
Remark 1.3.
$(Z_{m,n}\#(-Z_{m,n}),\tau)$ is a boundary of a cork which is not strong; that is, $\tau$ extends over some homology ball that $Z_{m,n}\#(-Z_{m,n})$ bounds. However, $\tau$ does not extend over any contractible 4-manifold that $Z_{m,n}\#(-Z_{m,n})$ bounds. See Remark 3.12.
@@
We prove Theorems 1.2, 1.4, 1.5, and 1.6 in Section 3, and we prove Proposition 1.7 in Section 4 (appendix).
@@ 3.1. Heegaard-Floer-theoretic methods
$h_{\iota\circ\tau}(Y)=[(CF^{-}(Y)[-2],\iota\circ\tau)]\in\mathfrak{I},$
where $\iota$ is the homotopy involution on $CF^{-}(Y)$ defined in [HM17]. The local equivalence group $\mathfrak{I}$ admits a partial order; see [DHM20, Definition 3.6]. The invariants $h_{\tau}(Y)$ and $h_{\iota\circ\tau}(Y)$ are monotonic under an appropriate equivariant negative-definite cobordism, as in the following theorem:
@@
We construct an interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)=(\Sigma(2,2m+1,4m+3),\sigma)$ as shown in Figure 13. This is formed by attaching (1)(-1)-framed 2-handles along the green unknots as in the top left of Figure 13. We see that the condition of Theorem 3.2 (2) are satisfied by simultaneously sliding the left green unknot over the right component of $Y_{m,1}$ and the right green unknot over the left component of $Y_{m,1}$. By Theorem 3.2 (2) and Lemma 3.3,
@@
Figure 13. The interchanging (1,1)(-1,-1)-cobordism from $(Y_{m,1},\sigma)$ to $(B_{m},\sigma)$ used in the proof of Theorem 1.4, obtained by attaching (1)(-1)-framed 2-handles along the green unknots.
@@
Remark 3.6.
It can be verified that $Y_{m,1}$ is diffeomorphic to the boundary $\partial W_{m}$, where $(W_{m},\tau)$ is the Akbulut-Yasui cork in [AY08]. In [DHM20, Theorem 1.13], it is shown that $h_{\tau}(\partial W_{m})\leq h_{\tau}(B_{m})<0$, which implies that $(\partial W_{m},\tau)$ is a strong cork.
Figure 14. The spinc-conjugating (1)(-1)-cobordism from $(Y_{m,n+1},\sigma)$ to $(Y_{m,n},\sigma)$ used in the proof of Theorem 1.4, obtained by attaching (1)(-1)-framed 2-handles along the green unknot
@@ 3.3. Instanton-theoretic methods
[ADMT23, Theorem 7.7] Let $\{(Y_{i},\tau_{i})\}^{\infty}_{i=1}$ be a sequence of oriented integer homology 3-spheres equipped with orientation-preserving involutions $\tau_{i}$. Assume that each $Y_{i}$ bounds a compact, contractible 4-manifold, and that the fixed-point set of each $\tau_{i}$ is a copy of $S^{1}$, so that the equivariant connected sum operation is well-defined. Suppose that:
@@
Figure 15. The equivariant negative-definite cobordism $W_{m,n}$ from $(Z_{m,n+1},\tau)$ to $(Z_{m,n},\tau)$ used in the proof of Theorem 1.2, obtained by attaching (1)(-1)-framed 2-handles along the green unknot.
@@
Remark 3.12.
$(Z_{m,n}\#(-Z_{m,n}),\tau)$ is not a strong cork, but $\tau$ does not extend over any contractible 4-manifold bounded by $Z_{m,n}\#(-Z_{m,n})$. If $\tau$ were to extend over some contractible 4-manifolds $W$ that $Z_{m,n}\#(-Z_{m,n})$ bounds, then we would obtain a simply-connected, equivariant definite cobordism $W^{\prime}$ from $(Z_{m,n},\tau)$ to itself. As in the proof of Theorem 1.2, $r_{0}(Z_{m,n},\tau)<\infty$. By Theorem 3.7, $r_{0}(Z_{m,n},\tau)<r_{0}(Z_{m,n},\tau)$, which is a contradiction.
@@
Figure 19. The simply-connected, equivariant negative-definite cobordism from $(S^{3}_{1/{m+1}}(K_{n}),\tau)$ to $(S^{3}_{1/m}(K_{n}),\tau)$ used in the proof of Theorem 1.6, obtained by attaching (1)(-1)-framed 2-handles along the green unknot.